OpenStudy (anonymous):

determine solution set: 4/5x+1=3/7 determine solution set: 4/5x+1=3/7 @Mathematics

jimthompson5910 (jim_thompson5910):

Does the problem look like \[\Large \frac{4}{5x+1}=\frac{3}{7}\] ???

OpenStudy (anonymous):

no --->